I'm confused as to what formula I would use for this word problem:
If the sides of a square are lengthened by 6cm, the area becomes 144cm^2. Find the length of the original square.

Answer by mananth(16946)

let the length of the square be x
increase by 6
= x+6
(x+6)^2=144
Take square root of the equation
x+6 = 12
x=12-6
x=6 cms the original length.